The odds against a particular candidate winning an election are estimated to be 5 to 8. If those odds are accurate, what is the probability that the candidate will win the election?
5
-
8 is the answer i came up with.

Answer by sudhanshu_kmr(1152)


odds against the probability = 5/8

probability that the candidate will win = 8/(5+8) = 8/13